# When a Reinforcement Learning Agent Encounters Unknown Unknowns

### v1

@@ Appendix I Optimism
$\leq|r_{h}(s,a)|+|\frac{1}{n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)[\mathring{p}_{h}^{k}\mathring{\overline{V}}_{h+1}^{k-1}(s,a)+\mathring{\gamma}_{h}^{k}(s,a)\mathring{p}_{h}^{k}(\mathring{\overline{V}}_{h+1}^{k-1}-\mathring{V}_{h,s,a}^{k-1})(s,a)]|$
$\leq 1+|\frac{1}{n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)[H+\frac{H^{2}}{2}|\leq H^{2}.$
@@
$\leq\frac{2}{n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)Var_{p_{h}}({V}_{h+1}^{\pi^{k}})(s,a)+\frac{2H}{n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)p_{h}(\mathring{\overline{V}}_{h+1}^{k-1}-V_{h+1}^{\pi^{k}})(s,a).$
@@
$\ \ \ \ \ +\frac{1}{Hlog(T)n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)p_{h}(\mathring{\overline{V}}_{h+1}^{k-1}-V_{h+1}^{\pi^{k}})(s,a)\frac{\zeta}{n_{h}^{t}(s,a)}+4H^{2}\frac{log(T)\zeta}{n_{h}^{t}(s,a)}.$ (8)
@@
$\beta_{h}^{t}(s,a)$ $\leq 2\sqrt{\frac{3}{\tilde{n}_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)Var_{p_{h}}(V_{h+1}^{\pi^{k}})(s,a)\frac{\zeta}{\tilde{n}_{h}^{t}(s,a)}}+106H^{3}\frac{log(T)\zeta}{\tilde{n}_{h}^{t}(s,a)}$
@@
$W_{h}^{t}(s,a)$ $\leq\frac{1}{n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a){p}_{h}(\mathring{\overline{V}}_{h+1}^{k-1})^{2}(s,a)-(\frac{1}{n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a){p}_{h}\mathring{\overline{V}}_{h+1}^{k-1}(s,a))^{2}$
@@
$=Var_{p_{h}}(V_{h+1}^{*})(s,a)+\frac{1}{n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a){p}_{h}((\mathring{\overline{V}}_{h+1}^{k-1})^{2}-(V_{h+1}^{*})^{2})(s,a)$
@@
$2\sqrt{W_{h}^{t}(s,a)\frac{\zeta}{n_{h}^{t}(s,a)}}$ $\leq 2\sqrt{\frac{3}{n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)Var_{p_{h}}(V_{h+1}^{\pi^{k}})(s,a)\frac{\zeta}{n_{h}^{t}(s,a)}}$
@@
$\leq 14H^{3}\frac{log(T)\zeta}{n_{h}^{t}(s,a)}+\frac{5}{4Hlog(T)n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)\mathring{\gamma}_{h}^{k}(s,a)p_{h}(\mathring{V}_{h,s,a}^{k-1}-\mathring{\overline{V}}_{h+1}^{k-1})(s,a)$
@@
We assume the event $\mathcal{D}$, Homeland condition in Section 4.2, and Assumption 1 hold.
From Lemma 6, we have that $\mathcal{D}$, Homeland condition in Section 4.2 and Assumption 1 hold with probability at least $1-\delta$. Fix $(s,a,h)\in\mathcal{S}\times\mathcal{A}\times[H]$ and $t\in[T]$.
@@
$\hat{R}_{h}^{T}(s,a)\triangleq\sum_{t=0}^{T-1}\chi_{h}^{t+1}(s,a)(\mathring{\overline{Q}}_{h}^{t}-Q_{h}^{\pi^{t+1}})(s,a)=\sum_{t=t(s)-1}^{T-1}\chi_{h}^{t+1}(s,a)(\mathring{\overline{Q}}_{h}^{t}-Q_{h}^{\pi^{t+1}})(s,a).$
@@
$\hat{R}_{h}^{T}(s,a)\leq$ $\sum_{t=t(s)-1}^{T-1}\chi_{h}^{t+1}(s,a)p_{h}(\mathring{V}^{t}_{h,s,a}-V_{h+1}^{\pi^{t+1}})(s,a)+130H^{3}\sum_{t=t(s)-1}^{T-1}\chi_{h}^{t+1}(s,a)\frac{log(T)\zeta}{\tilde{n}_{h}^{t}(s,a)}$
@@
$\hat{R}_{h}^{T}(s,a)\leq$ $44log(T)\sqrt{\zeta\sum_{t=t(s)-1}^{T-1}\chi_{h}^{t+1}(s,a)Var_{p_{h}}(V_{h+1}^{\pi^{t+1}})(s,a)}+1041H^{3}log(T)^{2}\zeta$
@@
$\sqrt{\sum_{t=t(s)-1}^{T-1}\chi_{h}^{t+1}(s,a)Var_{p_{h}}(V_{h+1}^{\pi^{t+1}})(s,a)}\leq\sqrt{2\sum_{t=t(s)-1}^{T-1}\overline{p}_{h}^{t+1}(s,a)Var_{p_{h}}(V_{h+1}^{\pi^{t+1}})(s,a)}+\sqrt{8\zeta}H$
@@
$\hat{R}_{h}^{T}(s,a)\leq$ $63log(T)\sqrt{\zeta\sum_{t=t(s)-1}^{T-1}\overline{p}_{h}^{t+1}(s,a)Var_{p_{h}}(V_{h+1}^{\pi^{t+1}})(s,a)}+1754H^{3}log(T)^{2}\zeta$
@@
$\hat{R}_{h}^{T}\triangleq\sum_{s\in S}\sum_{t=0}^{T-1}\overline{p}_{h}^{t+1}(s)(\mathring{\overline{V}}_{h}^{t}-V_{h}^{\pi^{t+1}})(s).$
We use the definition of $\mathring{\overline{V}}_{h}^{k}(s)$, Assumption 1 and the facts that $\mathcal{D}$ holds, $\mathring{\overline{Q}}_{h}^{k}\geq 0$ on $\mathcal{D}$ and that for all $x>1$ the following holds $\frac{1}{1-\frac{1}{4x}}\leq 1+\frac{1}{x}$ to get
$\ \ \ \ \ \sum_{t=0}^{T-1}\overline{p}_{h}^{t+1}(s)(\mathring{\overline{V}}_{h}^{t}-V_{h}^{\pi^{t+1}})(s)$
$=\sum_{t=0}^{(t(s)-2)\vee 0}\overline{p}_{h}^{t+1}(s)(\mathring{\overline{V}}_{h}^{t}-V_{h}^{\pi^{t+1}})(s)+\sum_{t=t(s)-1}^{T-1}\overline{p}_{h}^{t+1}(s)(\mathring{\overline{V}}_{h}^{t}-V_{h}^{\pi^{t+1}})(s)$
$\leq H(t(s)-1)+\frac{1}{1-\frac{1}{4H}}\sum_{t=t(s)-1}^{T-1}\chi_{h}^{t+1}(s)(\mathring{\overline{V}}_{h}^{t}-V_{h}^{\pi^{t+1}})(s)+\frac{56}{3}H^{2}\zeta$
$\leq H\sqrt{T}+(1+\frac{1}{H})\sum_{t=t(s)-1}^{T-1}\chi_{h}^{t+1}(s)(\mathring{\overline{V}}_{h}^{t}-V_{h}^{\pi^{t+1}})(s)+19H^{2}\zeta$
$\leq H\sqrt{T}+(1+\frac{1}{H})\sum_{t=t(s)-1}^{T-1}\chi_{h}^{t+1}(s)\pi_{h}^{t+1}(\mathring{\overline{Q}}_{h}^{t}-Q_{h}^{\pi^{t+1}})(s)+19H^{2}\zeta.$
@@
$\hat{R}_{h}^{T}$ $\leq HS\sqrt{T}+(1+\frac{1}{H})\sum_{s\in S}\sum_{t=t(s)-1}^{T-1}\chi_{h}^{t+1}(s)\pi_{h}^{t+1}(\mathring{\overline{Q}}_{h}^{t}-Q_{h}^{\pi^{t+1}})(s)+19H^{2}S\zeta$
$=HS\sqrt{T}+(1+\frac{1}{H})\sum_{(s,a)\in S\times A}\sum_{t=t(s)-1}^{T-1}\chi_{h}^{t+1}(s,a)(\mathring{\overline{Q}}_{h}^{t}-Q_{h}^{\pi^{t+1}})(s,a)+19H^{2}S\zeta$
$=HS\sqrt{T}+(1+\frac{1}{H})\sum_{(s,a)\in S\times A}\hat{R}_{h}^{T}(s,a)+19H^{2}S\zeta$
$\leq HS\sqrt{T}+126log(T)\sqrt{\zeta}\sum_{(s,a)\in S\times A}\sqrt{\sum_{t=t(s)-1}^{t-1}\overline{p}_{h}^{t+1}(s,a)Var_{p_{h}}(V_{h+1}^{\pi^{t+1}})(s,a)}$
$\ \ \ \ \ +(1+\frac{167}{H})\sum_{(s,a)\in S\times A}\sum_{t=t(s)-1}^{T-1}\overline{p}_{h}^{t+1}(s,a)p_{h}(\mathring{\overline{V}}_{h+1}^{t}-V_{h+1}^{\pi^{t+1}})(s,a)$
@@
$\ \ \ \ \ +126log(T)\sqrt{\zeta SA\sum_{(s,a)\in S\times A}\sum_{t=t(s)-1}^{t-1}\overline{p}_{h}^{t+1}(s,a)Var_{p_{h}}(V_{h+1}^{\pi^{t+1}})(s,a)}.$ (21)
@@
$\ \ \ \ \ \sum_{(s,a)\in S\times A}\sum_{t=t(s)-1}^{T-1}\overline{p}_{h}^{t+1}(s,a)p_{h}(\mathring{\overline{V}}_{h+1}^{t}-V_{h+1}^{\pi^{t+1}})(s,a)$
$=\sum_{(s,a,s^{\prime})\in S\times A\times S}\sum_{t=t(s)-1}^{T-1}\overline{p}_{h}^{t+1}(s,a)p_{h}(s^{\prime}|s,a)(\mathring{\overline{V}}_{h+1}^{t}-V_{h+1}^{\pi^{t+1}})(s^{\prime})$
@@
$\hat{R}_{1}^{T}$ $\leq\sum_{h=1}^{H}(1+\frac{167}{H})^{H-h}[126log(T)\sqrt{\zeta SA\sum_{(s,a)\in S\times A}\sum_{t=t(s)-1}^{t-1}\overline{p}_{h}^{t+1}(s,a)Var_{p_{h}}(V_{h+1}^{\pi^{t+1}})(s,a)}$
$\ \ \ \ \ +HS\sqrt{T}+3527SAH^{3}log(T)^{2}\zeta]$
@@
$\ \ \ \ \ +e^{167}H^{2}S\sqrt{T}+3527e^{167}SAH^{4}log(T)^{2}\zeta$
$\leq\tilde{\mathcal{O}}(\sqrt{\zeta H^{3}SAT})+\tilde{\mathcal{O}}(\zeta H^{4}SA)+\tilde{\mathcal{O}}(H^{2}S\sqrt{T}).$
@@
Appendix K Preparation Lemmas

### v2

@@ Appendix I Proof of Theorem 1
$\leq|r_{h}(s,a)|+$
$\ \ \ \ \left|\frac{1}{n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)\left[\mathring{p}_{h}^{k}\mathring{\overline{V}}_{h+1}^{k-1}(s,a)+\mathring{\gamma}_{h}^{k}(s,a)\mathring{p}_{h}^{k}(\mathring{\overline{V}}_{h+1}^{k-1}-\mathring{V}_{h,s,a}^{k-1})(s,a)\right]\right|$
$\leq 1+\left|\frac{1}{n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)\left[H+\frac{H^{2}}{2}\right]\right|\leq H^{2}.$
@@
$\leq\frac{2}{n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)Var_{p_{h}}({V}_{h+1}^{\pi^{k}})(s,a)$
$\ \ \ \ +\frac{2H}{n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)p_{h}(\mathring{\overline{V}}_{h+1}^{k-1}-V_{h+1}^{\pi^{k}})(s,a).$
@@
$\leq\sqrt{\frac{4}{n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)Var_{p_{h}}({V}_{h+1}^{\pi^{k}})(s,a)\frac{\zeta}{n_{h}^{t}(s,a)}}+4H^{2}\frac{log(T)\zeta}{n_{h}^{t}(s,a)}$
$\ \ \ \ +\frac{1}{Hlog(T)n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)p_{h}(\mathring{\overline{V}}_{h+1}^{k-1}-V_{h+1}^{\pi^{k}})(s,a)\frac{\zeta}{n_{h}^{t}(s,a)}.$ (13)
@@
$\beta_{h}^{t}(s,a)$ $\leq 2\sqrt{\frac{3}{\tilde{n}_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)Var_{p_{h}}(V_{h+1}^{\pi^{k}})(s,a)\frac{\zeta}{\tilde{n}_{h}^{t}(s,a)}}$
$\ \ \ \ +106H^{3}\frac{log(T)\zeta}{\tilde{n}_{h}^{t}(s,a)}$
@@
$W_{h}^{t}(s,a)$ $\leq\frac{1}{n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a){p}_{h}(\mathring{\overline{V}}_{h+1}^{k-1})^{2}(s,a)$
$\ \ \ \ -(\frac{1}{n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a){p}_{h}\mathring{\overline{V}}_{h+1}^{k-1}(s,a))^{2}$
@@
$=Var_{p_{h}}(V_{h+1}^{*})(s,a)$
$\ \ \ \ +\frac{1}{n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a){p}_{h}((\mathring{\overline{V}}_{h+1}^{k-1})^{2}-(V_{h+1}^{*})^{2})(s,a)$
@@
$\ \ \ \ 2\sqrt{W_{h}^{t}(s,a)\frac{\zeta}{n_{h}^{t}(s,a)}}$
@@
$\leq 14H^{3}\frac{log(T)\zeta}{n_{h}^{t}(s,a)}$
$\ \ \ \ +\frac{5}{4Hlog(T)n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)\mathring{\gamma}_{h}^{k}(s,a)p_{h}(\mathring{V}_{h,s,a}^{k-1}-\mathring{\overline{V}}_{h+1}^{k-1})(s,a)$
@@
We assume the event $\mathcal{D}$ and Homeland condition in Section 4.2.
From Lemma 6, we have that $\mathcal{D}$ and Homeland condition in Section 4.2 hold with probability at least $1-\delta$. Fix $(s,a,h)\in\mathcal{S}\times\mathcal{A}\times[H]$ and $t\in[T]$.
@@
$\hat{R}_{h}^{T}(s,a)$ $\triangleq\sum_{t=0}^{T-1}\chi_{h}^{t+1}(s,a)(\mathring{\overline{Q}}_{h}^{t}-Q_{h}^{\pi^{t+1}})(s,a)$
$=\sum_{t=t(s)-1}^{T-1}\chi_{h}^{t+1}(s,a)(\mathring{\overline{Q}}_{h}^{t}-Q_{h}^{\pi^{t+1}})(s,a).$
@@
$\ \ \ \ \hat{R}_{h}^{T}(s,a)$
$\leq\sum_{t=t(s)-1}^{T-1}\chi_{h}^{t+1}(s,a)p_{h}(\mathring{V}^{t}_{h,s,a}-V_{h+1}^{\pi^{t+1}})(s,a)+130H^{3}\sum_{t=t(s)-1}^{T-1}\chi_{h}^{t+1}(s,a)\frac{log(T)\zeta}{\tilde{n}_{h}^{t}(s,a)}$
@@
$\ \ \ \ \hat{R}_{h}^{T}(s,a)$
$\leq 44log(T)\sqrt{\zeta\sum_{t=t(s)-1}^{T-1}\chi_{h}^{t+1}(s,a)Var_{p_{h}}(V_{h+1}^{\pi^{t+1}})(s,a)}+1041H^{3}log(T)^{2}\zeta$
@@
$\ \ \ \ \ \sqrt{\sum_{t=t(s)-1}^{T-1}\chi_{h}^{t+1}(s,a)Var_{p_{h}}(V_{h+1}^{\pi^{t+1}})(s,a)}$
$\leq\sqrt{2\sum_{t=t(s)-1}^{T-1}\overline{p}_{h}^{t+1}(s,a)Var_{p_{h}}(V_{h+1}^{\pi^{t+1}})(s,a)}+\sqrt{8\zeta}H$
@@
$\hat{R}_{h}^{T}(s,a)\leq$ $63log(T)\sqrt{\zeta\sum_{t=t(s)-1}^{T-1}\overline{p}_{h}^{t+1}(s,a)Var_{p_{h}}(V_{h+1}^{\pi^{t+1}})(s,a)}$
$+1754H^{3}log(T)^{2}\zeta$
@@
$\hat{R}_{h}^{T}\triangleq\sum_{s\in S}\sum_{t=max\{t(s)-1,0\}}^{T-1}\overline{p}_{h}^{t+1}(s)(\mathring{\overline{V}}_{h}^{t}-V_{h}^{\pi^{t+1}})(s).$
We use the definition of $\mathring{\overline{V}}_{h}^{k}(s)$ and the facts that $\mathcal{D}$ holds, $\mathring{\overline{Q}}_{h}^{k}\geq 0$ on $\mathcal{D}$ and that for all $x>1$ the following holds $\frac{1}{1-\frac{1}{4x}}\leq 1+\frac{1}{x}$ to get
$\ \ \ \ \sum_{max\{t(s)-1,0\}}^{T-1}\overline{p}_{h}^{t+1}(s)(\mathring{\overline{V}}_{h}^{t}-V_{h}^{\pi^{t+1}})(s)$
$\leq\frac{1}{1-\frac{1}{4H}}\sum_{t=max\{t(s)-1,0\}}^{T-1}\chi_{h}^{t+1}(s)(\mathring{\overline{V}}_{h}^{t}-V_{h}^{\pi^{t+1}})(s)+\frac{56}{3}H^{2}\zeta$
$\leq(1+\frac{1}{H})\sum_{t=max\{t(s)-1,0\}}^{T-1}\chi_{h}^{t+1}(s)(\mathring{\overline{V}}_{h}^{t}-V_{h}^{\pi^{t+1}})(s)+19H^{2}\zeta$
$\leq(1+\frac{1}{H})\sum_{t=max\{t(s)-1,0\}}^{T-1}\chi_{h}^{t+1}(s)\pi_{h}^{t+1}(\mathring{\overline{Q}}_{h}^{t}-Q_{h}^{\pi^{t+1}})(s)+19H^{2}\zeta.$
@@
$\hat{R}_{h}^{T}$ $\leq(1+\frac{1}{H})\sum_{s\in S}\sum_{t=max\{t(s)-1,0\}}^{T-1}\chi_{h}^{t+1}(s)\pi_{h}^{t+1}(\mathring{\overline{Q}}_{h}^{t}-Q_{h}^{\pi^{t+1}})(s)+19H^{2}S\zeta$
$=(1+\frac{1}{H})\sum_{(s,a)\in S\times A}\sum_{t=max\{t(s)-1,0\}}^{T-1}\chi_{h}^{t+1}(s,a)(\mathring{\overline{Q}}_{h}^{t}-Q_{h}^{\pi^{t+1}})(s,a)+19H^{2}S\zeta$
$=(1+\frac{1}{H})\sum_{(s,a)\in S\times A}\hat{R}_{h}^{T}(s,a)+19H^{2}S\zeta$
$\leq 126log(T)\sqrt{\zeta}\sum_{(s,a)\in S\times A}\sqrt{\sum_{t=max\{t(s)-1,0\}}^{T-1}\overline{p}_{h}^{t+1}(s,a)Var_{p_{h}}(V_{h+1}^{\pi^{t+1}})(s,a)}$
$\ \ \ \ +(1+\frac{167}{H})\sum_{(s,a)\in S\times A}\sum_{t=max\{t(s)-1,0\}}^{T-1}\overline{p}_{h}^{t+1}(s,a)p_{h}(\mathring{\overline{V}}_{h+1}^{t}-V_{h+1}^{\pi^{t+1}})(s,a)$
@@
$\ \ \ \ +126log(T)\sqrt{\zeta SA\sum_{(s,a)\in S\times A}\sum_{t=max\{t(s)-1,0\}}^{t-1}\overline{p}_{h}^{t+1}(s,a)Var_{p_{h}}(V_{h+1}^{\pi^{t+1}})(s,a)}.$ (26)
@@
$\ \ \ \ \sum_{(s,a)\in S\times A}\sum_{t=max\{t(s)-1,0\}}^{T-1}\overline{p}_{h}^{t+1}(s,a)p_{h}(\mathring{\overline{V}}_{h+1}^{t}-V_{h+1}^{\pi^{t+1}})(s,a)$
$=\sum_{(s,a,s^{\prime})\in S\times A\times S}\sum_{t=max\{t(s)-1,0\}}^{T-1}\overline{p}_{h}^{t+1}(s,a)p_{h}(s^{\prime}|s,a)(\mathring{\overline{V}}_{h+1}^{t}-V_{h+1}^{\pi^{t+1}})(s^{\prime})$
@@
$\hat{R}_{1}^{T}$ $\leq\sum_{h=1}^{H}(1+\frac{167}{H})^{H-h}$
$\ \ \ \ \cdot[126log(T)\sqrt{\zeta SA\sum_{(s,a)\in S\times A}\sum_{t=t(s)-1}^{t-1}\overline{p}_{h}^{t+1}(s,a)Var_{p_{h}}(V_{h+1}^{\pi^{t+1}})(s,a)}$
$\ \ \ \ +3527SAH^{3}log(T)^{2}\zeta]$
@@
$\ \ \ \ +3527e^{167}SAH^{4}log(T)^{2}\zeta$
$\leq\tilde{\mathcal{O}}(\sqrt{\zeta H^{3}SAT})+\tilde{\mathcal{O}}(\zeta H^{4}SA).$
@@
Appendix J Preparation Lemmas